# Empowering Scientific Workflows with Federated Agents

### v2

@@ 4.2.1. Data Transfer
We first investigate the pass-by-reference and direct communication optimizations of the distributed exchange. In baseline, all message data are communicated indirectly between peers via the exchange’s object store. The object store is located remotely on a Chameleon Cloud node. In pass-by-ref, messages are still communicated with the object store, but action arguments and results are replaced with references using ProxyStore. ProxyStore is configured to use ZeroMQ and ProxyStore’s P2P endpoints for intra-site and inter-site transfer of referenced objects, respectively. In direct, messages are communicated directly between peers, circumventing the cloud-hosted object store; this is only possible when peers are located within the same site.
In Figure 4 (top), we measure the time it takes for a client to invoke a no-op action on an agent as a function of input and output payload size. We compare baseline, pass-by-ref, and direct across two scenarios: Aurora → Aurora, where the client and the agent are located on two different Aurora nodes, and Workstation → Aurora, where the client is located on a personal workstation and the agent is located on an Aurora node. The latencies between the three sites are Aurora to Chameleon: 31 ms; Aurora to Workstation: 12 ms; and Workstation to Chameleon: 42 ms. The workstation is limited to an 800 Mbps internet connection.
We observe that network latency to the exchange object store limits performance at smaller payload sizes ($\leq$ 100 KB). Direct, which is possible only in the intra-site scenario, circumvents these latencies. In both scenarios, pass-by-ref alleviates overheads of data transfer to and from the object store by communicating data directly between the client and agent via ProxyStore. For intra-site transfers, pass-by-ref and direct reduce action latency compared to the baseline by 91.2% and 97.6%, respectively, with 100 MB payloads. For inter-site transfers, pass-by-ref reduces action latency by 78.8%.
Pass-by-ref also reduces overheads when actions pass data to subsequent actions, a common pattern in multi-agent systems. We evaluated this optimization by measuring the round-trip time of action chains in which data are passed through $n$ actions, each invoked on a separate agent, and results are returned through each agent as well. Pass-by-ref reduces the size of messages communicated via the exchange, as indicated by the shallower slope in Figure 4 (bottom). Data are only communicated once to the agent that uses the data (here, the last agent in the chain).
4.2.2. Handle Multiplexing
Figure 5. Maximum no-op action throughput for a single agent requesting actions from $n$ worker agents. The handle multiplexing optimization improves performance by reducing the number of mailbox listener threads from $n$ to 1.
@@ 4.2.2. Handle Multiplexing
[Handle multiplexing ablation study.]
As described in Section 3.3.3, the communication of multiple handles within a process is multiplexed through a single mailbox. Without this optimization, each handle held by a client process or agent would create a thread for communication. We evaluated this optimization in Figure 5 by creating one agent that submits a bag-of-tasks to $n$ worker agents and comparing the task throughput with (multiplex) and without (baseline) mailbox multiplexing. Multiplexing improves throughput by 41.7% with 52 worker agents due to reduced threading overheads.
4.3. Agent Messaging
Here, we investigate the performance of agent messaging. As in Section 4.1, the object store of the exchange is located on the head node of the Aurora batch job.
@@ 4.3. Agent Messaging
4.3.1. Action Latency
In Figure 6 (top), we show action latency—the time between sending an action request and receiving a result— between two agents on different nodes. We vary the input/output payload size to understand data transfer overheads. The mean and standard deviation roundtrip latencies are 385$\pm$301 $\mu$s in Academy, 1186$\pm$1059 $\mu$s in Dask, and 526$\pm$308 $\mu$s in Ray for the smallest 10 KB payloads, with latencies increasing with payload size.
4.3.2. Action Throughput
We measure the maximum action throughput for a single agent by submitting a bag of no-op tasks to a pool of worker agents (following the “agents as process pools” pattern from Section 3.4.4). The pool contains 208 agents across four nodes to ensure that each worker agent is not over-saturated with work. That is, the single submitter agent is the limiting factor for performance. Academy, Dask, and Ray achieve maximum throughputs of 3.4K, 185, and 14.1K action/s, respectively. Academy is 18$\times$ faster than Dask but 4$\times$ slower than Ray; however, this is a worst case scenario with no-op tasks and we believe >3K actions/s to be sufficient for real-world agents.
4.3.3. Agent Conversations
@@
In Figure 6 (bottom), we simulate a common pattern in LLM agents where two agents have a back-and-forth conversation. We compare Academy to AutoGen, a popular framework for creating multi-agent AI applications. Each agent is run in a different process on the same node. Agents send ten messages back-and-forth, and we repeat with varying message sizes to simulate different kinds of conversations (i.e., text-only versus multi-modal). AutoGen’s distributed agent runtime uses gRPC which has a maximum message size of 4 MB. Academy has comparatively lower overhead messaging in distributed settings.
4.4. Memory Overhead
@@
Figure 7. Memory used by $n$ agents/actors. We encountered Ray crashes when deploying 104 actors on a single Aurora node (i.e., all cores on both sockets).
@@
[Memory used by $n$ actors/agents.]
We show memory used as a function of number of agents in Figure 7; for Academy, we compare two launchers: a low-overhead but single-node process-pool and Parsl’s High-throughput Executor. For fairness, we disable features in Dask and Ray that may reduce performance, such as dashboards, and set the initial Ray object store size to the smallest possible value. Academy agents have low memory overheads, making them suitable for memory-constrained devices, such as when deployed across edge devices via the Globus Compute launcher. The Ray cluster head worker has high memory overhead, but that initial overhead is amortized as the number of actors is increased, indicating that each actor has modest overhead.
5. Case Studies
We use three applications to demonstrate the practicality, generality, and robustness of our system in real-world settings. These examples illustrate how Academy integrates with existing research infrastructure, supports a range of computational patterns, and adapts to the varying demands of scientific applications. They validate key design choices, uncover integration challenges, and provide guidance to researchers building agentic workflows.
5.1. Materials Discovery
Figure 8. We execute MOFA by deploying agents across federated infrastructure with Globus Compute. The Assembler, Database, and Estimator run on Chameleon Cloud nodes with fast single-core performance; the Generator and Validator run on Aurora login nodes and execute AI and simulation tasks, respectively, on Aurora compute nodes; and the Optimizer runs on a Polaris login node and executes simulation tasks on Polaris compute nodes. Each agent is responsible for a single MOFA stage, and agents cooperate through message passing, such as to request more work and trigger periodic events. The agents on Aurora and Polaris use Parl to scale resources up and down based on workload needs.
[…]
[Agentic MOFA workflow design.]
MOFA (Yan et al., 2025) is an online learning application for generating, screening, and evaluating metal organic frameworks (MOFs) that couples generative AI methods with computational chemistry. MOFs are polymers composed of inorganic metal clusters and organic ligands; their porosity and large surface area make them suitable for gas adsorption applications such as carbon capture (Furukawa et al., 2013). The goal of MOFA is to generate high-performing candidates by intelligently navigating the intractable combinatorial space of possible MOF structures. MOFA is representative of a broad class of scientific workflows that require careful integration of heterogeneous tasks spanning AI and simulation.
MOFA involves five stages: (1) a generative AI model produces candidate ligands; (2) these ligands are combined with predefined metal clusters to assemble candidate MOFs; (3) the candidates undergo iterative screening and validation using a series of molecular dynamics simulations; (4) CO2 adsorption properties of the most promising structures are simulated and recorded in a database; and (5) the generative model is periodically retrained on the accumulated results to enhance its performance over time. MOFA utilizes Colmena (Ward et al., 2021) to coordinate the flow of data between stages and to distribute computations across CPU and GPU resources within a single batch job. However, this design has key limitations: stages cannot be deployed across heterogeneous resources, such as to leverage hardware best optimized for the specific computations; stages cannot independently scale in or out—resources are bound by the size of a single job; integrating new components within tightly coupled code is challenging; and integration with asynchronous processes, such as synthesis in a real laboratory, are infeasible.
Figure 9. Execution trace of the agentic MOFA workflow of Figure 8 over three hours. (Top) Active tasks per agent. The vertical axis height represents the maximum size of the resource pool allocated by each agent (i.e., CPUs or GPUs). Assembler tasks are short and infrequent. (Middle) Cumulative tasks submitted per agent. (Bottom) Active workers allocated in each agent’s resource pool. Worker allocations vary with demand (as in Assembler and Estimator) or batch job wall times (as in Generator, Validator, and Optimizer).
[…]
[Agentic MOFA workflow execution trace.]
@@ 5.1. Materials Discovery
MOFA is an excellent candidate for an agentic workflow, as we demonstrate by porting MOFA to use Academy and deploying the workflow across federated resources: see Figure 8. We express MOFA through six agents: Database, Generator, Assembler, Validator, Optimizer, and Estimator. Each agent is responsible for a different component of the workflow and manages its own resources (i.e., storage and compute). Agents are remotely deployed across Chameleon Cloud nodes and the login nodes of Aurora and Polaris via Globus Compute, and communicate via the distributed exchange backed by a Redis instance in Chameleon Cloud.
@@
An execution trace of the agentic MOFA workflow (Figure 9) shows how each agent scales out its allocated resources as work becomes available, and in the case of Assembler and Estimator, scale down when their workload decreases. The Generator, Validator, and Optimizer consistently have work to do but their batch jobs within which workers run have 60 minute wall times that expire and then must be resubmitted, causing temporary drops the the number of workers. Active tasks that are killed are automatically restarted in the next job. This separation of concerns is key to enabling long-running workflows—resource infrastructure is not persistently available and agents will need to be able to adapt to that varying availability. A second benefit of this model is the loose coupling between agents. For example, the specific implementation of a given agent can be trivially swapped provided the behavior (i.e., the API that agents expose) remains the same. In addition, it becomes easier to integrate future agents, such as to incorporate embodied agents that interact with self-driving labs to synthesize and evaluate the best-performing MOFs in the real-world. While automated MOF synthesis is not yet practical, the capabilities of self-driving labs are rapidly improving (Abolhasani and Kumacheva, 2023; Vescovi et al., 2023), and it is tangible to envision a future where these loosely coupled agentic workflows incorporate services provided by self-driving labs through embodied agents.
5.2. Decentralized Learning
@@
In decentralized machine learning a set of models learn collaboratively across distributed datasets (Hegedűs et al., 2019). This paradigm is particularly relevant today as data are increasingly generated in decentralized settings and transfer to a centralized location can be infeasible for cost and privacy reasons. Each device in a decentralized learning workflow performs three steps: (1) train a model on local data for a set number of iterations; (2) receive models from neighboring devices and send its own model to neighbors; and (3) update the local model via an all reduce operation performed across its own and received models. Reframing the decentralized learning workflow as an agentic workflow is a natural and powerful extension.
@@
We implement a decentralized and asynchronous machine learning exemplar using Academy. The agents and the communication channels between them can be represented as a graph where nodes are agents and edges are communication channels. We choose a powerlaw cluster graph to approximate real-world networks (Holme and Kim, 2002). Each agent is responsible for training its local model, receiving neighboring agents’ models, and aggregating received models with its own model on a periodic basis. Each agent uses a copy of the MNIST dataset (Deng, 2012). The agents are configured to use pass-by-ref with ProxyStore as the transfer backend. Thus data communication between agents follows the network topology. We investigate the cost of distributing updates from all agents as we scale the size of the graph for different model sizes in Figure 10. We do not show training and aggregation time as it is approximately the same for all model sizes and does not increase with the number of agents. The agents are deployed on Aurora using Parsl, where each agent is pinned to a single GPU tile (two tiles per physical GPU), allowing 12 agents per node. Our results demonstrate Academy’s ability to support more than 1500 autonomous agents working collaboratively with no client coordination (as can be seen by the constant time in Figure 10).
@@
Figure 10. Model communication time to an agent’s neighbors averaged over five rounds of decentralized training. Training time and aggregation time are excluded since they are nearly constant.
[…]
5.3. Information Extraction
@@
Figure 11. Execution trace of the agentic MCQ workflow processing 10 manuscripts to generate and validate questions and answers over 15 minutes. The figure shows the active agents and the duration of their tasks. Agents employ either the Mistral-7B-Instruct-v0.3 or Meta-Llama-3-70B-Instruct model, denoted A and B, respectively.
[…]
@@
Exponential growth in scientific publications (Bornmann et al., 2021) creates potential for cross-disciplinary insights that are largely untapped due to the limitations of manual literature review. Automating information extraction from this vast and varied body of work using AI is crucial to accelerate scientific progress. AI methods can be employed to identify and synthesize key findings, methodologies, and datasets across fields and thus to identify connections and facilitate the cross-pollination of ideas that would otherwise go unnoticed (Buchkremer et al., 2019; Sourati and Evans, 2023).
@@
To explore the potential of agentic workflows for thus analyzing the scientific literature we used Academy to implement a system for generating and validating multi-choice questions (MCQs) from research publications (Ch and Saha, 2018; Catlett and Foster, 2025). The workflow includes a PDFParser agent to extract text from a manuscript; two Generator agents that use different LLMs to generate MCQs; an MCQSelector to choose subsets of questions to evaluate; and two MCQAnswerers and two AnswerScorers (again, each with a different LLM) to generate and validate, respectively, answers to questions. The agents use the Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) and Meta-Llama-3-70B-Instruct (at Meta, 2024) models, denoted A and B, respectively.
@@
6. Related Work

### v3

@@ IV-C1 Action Latency
IV-C2 Action Throughput
@@
IV-C3 Agent Conversations
@@
In Figure 5 (bottom), we simulate a common pattern in LLM agents where two agents have a back-and-forth conversation. We compare Academy to AutoGen, a popular framework for creating multi-agent AI applications. Each agent is run in a different process on the same node. Agents send 10 messages back and forth, repeating with varying message sizes to simulate different kinds of conversations (i.e., text-only vs. multi-modal). AutoGen’s distributed agent runtime uses gRPC which has a maximum message size of 4 MB. Academy has comparatively lower overhead messaging in distributed settings.
IV-D Memory Overhead
@@
Figure 6: Memory used by $n$ agents/actors. We encountered a bug causing Ray to crash when deploying 104 actors on a single Aurora node (i.e., all cores on both sockets).
We show memory used as a function of number of agents in Figure 6; for Academy, we compare two executors: a low-overhead but single-node process-pool and Parsl’s High-throughput Executor. For fairness, we disable features in Dask and Ray that may increase memory, such as dashboards, and set the initial Ray object store size to the smallest possible value. Academy agents have low memory overheads, making them suitable for memory-constrained devices, as when deployed to edge devices via Globus Compute. The Ray cluster head worker has high memory overhead, but that initial overhead is amortized as the number of actors is increased, indicating that each actor has modest overhead.
V Case Studies
We use four applications to demonstrate the practicality, generality, and robustness of Academy in real-world settings. These examples illustrate how Academy integrates with existing research infrastructure, enables distinct capabilities, and adapts to the varying demands of scientific applications. More generally, these examples demonstrate how the agent paradigm simplifies the construction of the scientific application patterns identified in Section II.
V-A Materials Discovery
Figure 7: MOFA deploys agents across federated infrastructure with Globus Compute. The Assembler, Database, and Estimator run on Chameleon Cloud nodes with fast single-core performance; the Generator and Validator run on Aurora login nodes and execute AI and simulation tasks, respectively, on Aurora compute nodes; and the Optimizer runs on a Polaris login node and executes simulation tasks on Polaris compute nodes. Each agent is responsible for a single MOFA stage. Agents cooperate by passing messages on the exchange, such as to request more work and trigger periodic events. Agents on Aurora and Polaris use Parsl to scale resources up and down based on workload needs.
MOFA [24] is an online learning application for generating, screening, and evaluating metal organic frameworks (MOFs) that couples generative AI methods with computational chemistry. MOFs are polymers composed of inorganic metal clusters and organic ligands that are particularly suitable for gas adsorption applications such as carbon capture [40]. The goal of MOFA is to generate high-performing candidates by intelligently navigating space of possible MOF structures. MOFA is representative of a broad class of scientific workflows that require careful integration of heterogeneous tasks spanning AI and simulation.
MOFA involves five stages: (1) a generative AI model produces candidate ligands; (2) these ligands are combined with predefined metal clusters to assemble candidate MOFs; (3) the candidates undergo iterative screening and validation using a series of molecular dynamics simulations; (4) CO2 adsorption properties of the most promising structures are simulated and recorded in a database; and (5) the generative model is periodically retrained on the accumulated results to enhance its performance over time.
TABLE I: Per-node throughputs, in MOFs/min, for the various components of the MOFA application (Figure 7).
Task Academy Colmena/Parsl
Machine Throughput Machine Throughput
Validate Aurora 10.73 Polaris 4.34
Optimize Polaris 0.85 Polaris 0.85
Estimate Cloud 6.81 Polaris 1.85
@@ V-A Materials Discovery
We adapt MOFA to use Academy and deploy the resulting agentic application across federated resources: see Figure 7. We use six agents: Database, Generator, Assembler, Validator, Optimizer, and Estimator. Each agent is responsible for a different component of the workflow and manages its own resources (i.e., storage and compute). Agents are remotely deployed across Chameleon Cloud, Aurora and Polaris.
Figure 8: Execution trace of the Academy MOFA workflow. (Top) Active tasks per agent. The vertical axis height represents the maximum size of the resource pool allocated by each agent. (Middle) Cumulative tasks submitted per agent. (Bottom) Active workers allocated in each agent’s resource pool. Worker allocations vary with demand (as in Assembler and Estimator) or batch job wall times (as in Generator, Validator, and Optimizer).
@@
An execution trace of the agentic MOFA workflow (Figure 8) shows how each agent scales out its allocated resources as work becomes available, and in the case of the Generator, Validator, and Assembler, scale down when their workload decreases. After a ramp up period, the Optimizer consistently has work to do but their batch jobs within which workers run have 60 minute wall times that expire and then must be resubmitted, causing temporary drops in the number of workers. Active tasks that are killed are restarted in the next job. This separation of concerns is key to enabling long-running workflows—resource infrastructure is not persistently available and agents will need to be able to adapt to that varying availability.
We contrast the Academy implementation of MOFA to a version implemented with Colmena [21], a bespoke simulation campaign framework built on Parsl [32]—a traditional workflow system that does not support stateful federated agents. We compare the Academy version, deployed on federated resources, to the Colmena version run as a single batch job on Polaris. In addition to using different resources, the Academy execution includes queue wait times to acquire resources while the Colmena execution does not. Although the performances are thus not directly comparable, the comparison against Colmena/Parsl (shown in Figure 9) demonstrates how federated agents support more efficient resource utilization. With Academy, resource allocations (Figure 8) are managed autonomously by each agent, whereas Colmena allocates resources statically for the entire workflow. This, along with the hardware configuration of Polaris, means that when the Colmena workflow performs CPU-only tasks like “estimate,” the GPU resources of the allocation sit idle.
Figure 9: Resources used for a traditional workflow implementation of MOFA deployed on Polaris. The static resource configuration does not match the workload, so in the estimation stage, when tasks are only using the CPU, the rest of the resources are idle.
A further advantage is that the loose coupling between agents in the Academy model makes it trivial to swap one agent implementation for another, provided that the API exposed by the agent remains the same. It also becomes easier to integrate future agents, such as to incorporate embodied agents that interact with self-driving labs to synthesize the best-performing MOFs via physical experiments [4, 28, 41].
V-B Astronomical Spectroscopy
Integral field spectroscopy is a powerful technique used in observational astrophysics to gather spatially-resolved spectroscopic data from a target field at once, enabling the study of spatially-complex objects like distant strongly-lensed galaxies—optimal candidates for understanding galaxy formation in the early universe [42]. Before analyzing the spectra offered by the Integral Field Units for Magellan (IFU-M) instrument, scientists first process and calibrate the data which involves the overarching steps: pre-processing, spatial and wavelength calibrations, and datacube creation. A crucial step in pre-processing the data is detecting and filtering cosmic rays through subtracting multiple images of the same patch of sky. However, the optics of the telescope shift over time in response to small variations in external factors such as temperature. The images must be aligned by finding offset and scaling parameters that minimize the variation between images. Previously, the alignment was found through a grid search of parameters between pairs of images.
We employ a hybrid pattern for this workflow in which Academy conceptually maps the instrument into the workflow by estimating and storing the optical parameters of the telescope (see Figure 10). As these parameters vary smoothly in time, the agent uses previously calibrated images to inform the search for new parameters, leading to faster alignment and reduced noise in the processed images. The two spectrographs of the instrument are each represented by their own agent. Other tasks in the application remain stateless, allowing them to be scheduled to any available worker. In the future, deploying the agents on the instrument would allow the agent to use observations (e.g., temperature measurements) to inform the alignment, and provide on-site feedback for targeting observations.
Figure 10: Execution trace of the IFU-M workflow. Each color corresponds to a different task type. Tasks within the Parsl workflow are stateless, but invoke actions on agents that track the instrument state.
V-C Decentralized Learning
@@
In decentralized machine learning a set of models learn collaboratively across distributed datasets [43]. This paradigm is particularly relevant today as data are generated in decentralized settings and transfer to a centralized location can be infeasible for cost and privacy reasons. Each device in a decentralized learning workflow performs three steps: (1) train model on local data for a set number of iterations; (2) receive models from neighboring devices and send its own model to neighbors; and (3) update the local model via an all-reduce operation performed across its own and received models. It is straightforward to reframe such a decentralized learning workflow as an agentic workflow. We represent the application as a graph in which nodes are agents and edges are communication channels. Each agent trains its local model, receives neighboring agents’ models, and periodically aggregates received models with its own model.
@@
Using Academy, we simulate decentralized learning on Aurora. For the connectivity between devices, we choose a power-law cluster graph to approximate real-world networks [44]. Each agent uses a copy of the MNIST dataset [45]. The agents are configured to use pass-by-ref with ProxyStore as the transfer backend. We investigate the cost of distributing updates from all agents as we scale the size of the graph for different model sizes: see Figure 11. The agents are deployed on Aurora using Parsl, with each agent pinned to a single GPU tile (two tiles per physical GPU), allowing 12 agents per node. Our results demonstrate Academy’s ability to support more than 1500 autonomous agents working collaboratively with no client coordination, as shown by the constant times seen in Figure 11.
@@
Figure 11: Model communication time to an agent’s neighbors averaged over five rounds of decentralized training. Training time and aggregation time are excluded since they are nearly constant.
V-D Information Extraction
@@
Figure 12: Execution trace of the agentic MCQ workflow processing 10 manuscripts to generate and validate questions and answers over 15 minutes. The figure shows the active agents and the duration of their tasks. Agents employ either the Mistral-7B-Instruct-v0.3 or Meta-Llama-3-70B-Instruct model, denoted A and B, respectively.
@@
Exponential growth in scientific publications [46] creates potential for cross-disciplinary insights that are largely untapped due to the limitations of manual literature review. Automating information extraction from this vast and varied body of work using AI is crucial to accelerate scientific progress. AI methods can be employed to identify and synthesize key findings, methodologies, and datasets across fields and thus to identify connections and facilitate the cross-pollination of ideas that would otherwise go unnoticed [47, 48].
@@
To explore the potential of agentic workflows for thus analyzing the scientific literature, we used Academyto implement a system for generating and validating multi-choice questions (MCQs) from research publications [49, 50]. The workflow includes a PDFParser agent to extract text from a manuscript; two Generator agents that use different LLMs to generate MCQs; an MCQSelector to choose subsets of questions to evaluate; and two MCQAnswerers and two AnswerScorers (again, each with a different LLM) to generate and validate, respectively, answers to questions. The agents use the Mistral-7B-Instruct-v0.3 [51] and Meta-Llama-3-70B-Instruct [52] models, denoted A and B, respectively.
@@
VI Discussion
We describe several lessons learned implementing Academy and applying it to a diverse range of applications.
Scientific agentic systems require federation and statefulness. Science is often collaborative and multi-institutional. Research infrastructure (e.g., instruments, HPC systems, etc.) mirrors this inherently federated ecosystem. Building intelligent scientific systems requires agents that run in multiple locations and collaborate across institutions. Existing workflow systems (e.g., Parsl, Ray, Dask, Globus Compute) either do not support the capabilities necessary for agentic science, or face significant barriers when deployed on research infrastructure. Similarly, conversational agentic frameworks (e.g., LangChain, PydanticAI) are not designed for such environments, but can leverage Academy to deliver these capabilities.
Start-up barriers hinder adoption and deployment. Academy initially required that users deploy the hybrid exchange to allow inter-site communication. This requirement hindered the transition from local development to multi-site deployment. The introduction of the cloud exchange reduced barriers and thus enabled significant improvements (e.g., agent sharing and reuse). It also simplified integration with existing execution and data transfer solutions (e.g., Globus Compute, ProxyStore), further reducing barriers of using Academy.
Agentic middleware should not bias the behaviors that users create. Commonly used agentic systems such as LangChain and SmolAgents shoehorn applications into specific patterns (e.g., LLMs with tool-calling). Scientific agents behave in diverse ways that often do not fit such patterns. Rather than design specific behaviors into the system, we implemented patterns in Academy (e.g., agent hierarchies, agents as resource managers) and reused them in different contexts.
LLMs combined with tool calling cannot currently construct complex real-world science applications. For instance, in the MOFA workflow we experimented with the use of LLMs for such functions as improving MOF design and deciding which simulators and simulations to run. However, we found that the large number of tools and tool parameters, often with limited documentation, and sometimes subtle differences among tools, led to frequent problems including improperly configured/invoked codes that ran without errors but produced incorrect results. Manual construction of both tools (as Academy actions) and workflow architecture (by invoking handles) informed by domain expertise, created more reliable applications.
Autonomy and adaptability can improve application performance and resource utilization. Static resource allocations often inadequately match the dynamic nature of scientific workloads. Agentic systems enable not only component autonomy but also dynamic resource utilization at fine granularity.
VII Related Work